OpenStudy (anonymous):

If f(x) = 2x - 5, then f(4) is _____. 17 5 4 3

OpenStudy (anonymous):

f(4)=2(4)-5 f(4)=8-5 f(4)=3

OpenStudy (anonymous):

thanks